OpenStudy (anonymous):

Help simplifying the 5th root of a number. \[\sqrt[5]{125x^{15}y^{4}z^{25}}\]

OpenStudy (anonymous):

\[(x ^{15})^{1/5} = x ^{3}\]

OpenStudy (anonymous):

thats a hint right there!

OpenStudy (anonymous):

I know what to do with the multiples of 5 in the problem, my question is really what do you do with that 4?

Parth (parthkohli):

\[\sqrt[5]{125} \times \sqrt[5]{x^{15}} \times \sqrt[5]{y^4} \times \sqrt[5]{z^{25}}\]

OpenStudy (anonymous):

Okay, thanks. That helps a lot!

Parth (parthkohli):

You can simplify that further... :-)